# RevCD - Reversed Conditional Diffusion for Generalized Zero-Shot Learning

### v1

@@ Abstract
In Generalized Zero-Shot Learning (GZSL), we aim to recognize both seen and unseen categories using a model trained only on seen categories. In computer vision, this translates into a classification problem, where knowledge from seen categories is transferred to unseen categories by exploiting the relationships between visual features and available semantic information, such as text corpora or manual annotations. However, learning this joint distribution is costly and requires one-to-one training with corresponding semantic information. We present a reversed conditional Diffusion-based model (RevCD) that mitigates this issue by generating semantic features synthesized from visual inputs by leveraging Diffusion models’ conditional mechanisms. Our RevCD model consists of a cross Hadamard-Addition embedding of a sinusoidal time schedule and a multi-headed visual transformer for attention-guided embeddings. The proposed approach introduces three key innovations. First, we reverse the process of generating semantic space based on visual data, introducing a novel loss function that facilitates more efficient knowledge transfer. Second, we apply Diffusion models to zero-shot learning—a novel approach that exploits their strengths in capturing data complexity. Third, we demonstrate our model’s performance through a comprehensive cross-dataset evaluation. The complete code will be available on GitHub.
Keywords Zero-Shot learning  $\cdot$ Diffusion Model  $\cdot$ Generative ZLS  $\cdot$ Transfer Learning
@@ 1 Introduction
Zero-shot learning (ZSL) represents the state-of-the-art advancement in the domains of machine learning transferability and computer vision classification. By pushing the boundaries of knowledge extraction, we enable ML models to expand without costly retraining. This learning paradigm is particularly crucial as it addresses the inherent limitation of traditional machine learning models that require prior access to expensive datasets. ZSL leverages auxiliary knowledge, allowing models to explore unobserved events, edge cases, or new compositions without any additional training. Traditional approaches in ZSL focused on aligning attributes directly with object categories [26], while deep learning’s potential to create a joint embedding space of visual and semantic features [8, 35, 7, 12, 41] have rendered this approach obsolete. The shift towards latent-based methods highlights the importance of embedding space techniques because of their ability to decode and infer complex data distributions. This offers a promising resolution to the two main challenges of ZSL, the semantic gap, and generalization abilities [44].
@@
Our contribution introduces a Diffusion-based generative model, a notable innovation in ZSL. Distinct from conventional models that predominantly rely on attribute matching or embedding strategies, our RevCD model utilizes a diffusion process to model the data distributions iteratively, see Fig. 1. This augments the model’s capability to manage class variability and enhances its generalization capacity. This control over the latent space is required for overcoming the challenges of bias and hubness commonly encountered in ZSL methodologies [27].
@@ 2 Related works
Advancements in likelihood-based models have been central to the progress of zero-shot learning [37]. By framing the learning process as a maximum likelihood estimation problem, these methods effectively model data distributions, allowing for robust generalization across both seen and unseen classes. This section categorizes ZSL approaches according to the foundational models employed for latent space approximation, including Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Hybrid models. Additionally, it highlights the role of attention mechanisms and embedding strategies in enhancing these models’ performance.
VAE-based. Variational Autoencoders [25] have played a crucial role in ZSL due to their probabilistic framework for modeling latent spaces. Their adaptability in synthesizing unseen class prototypes, as demonstrated by [39], [5], and [23], underscores their versatility in ZSL applications. For instance, [18] incorporates a semantic-guided approach within a VAE framework, while [46] employs a decoupling strategy to enhance performance. A significant advantage of VAEs lies in their capacity to explicitly approximate data density. However, a key limitation is their tendency to generate blurry or overly smoothed features due to posterior collapse [33], which can obscure essential class-specific details needed to distinguish between unseen classes.
GAN-based. Generative Adversarial Networks [14] offer a powerful and dynamic framework for feature synthesis. GANs have been successfully adapted to generate features for unseen classes; [13], [21], [15], and [53]. GANs ability to produce sharp and realistic features through implicit density estimation makes them particularly effective for capturing fine-grained details. However, they are also prone to challenges such as training instability and mode collapse [4], which can result in a lack of diversity in the generated features. This limitation may hinder the model’s ability to accurately represent the full spectrum of unseen classes.
Hybrids. Hybrid models in ZSL leverage multiple architectures to enhance performance. The majority of hybrid frameworks incorporate sequential modules, as seen in [16], [10] and [32]. Employing a VAE to learn an embedding function that constrains the semantic or visual space allows for greater control over the generation of synthesized features. Nevertheless, the complexity and computational overhead of combining multiple models can pose challenges, especially in terms of model interpretability and scalability.
@@
Our contribution introduces a reversed Diffusion-based model (RevCD) for zero-shot inference. Diffusion models have been applied to improve accuracy as generative classifiers [6, 3, 40], and their capacity to generate synthetic data has been used to classify unseen compositions [9, 28]. However, these are limited by pre-training on prompt categories. Their implementation in a pure zero-shot setting is still absent. We address these limitations by leveraging the reversed process for generating conditioned semantic embeddings, aiming for effective generalization to unseen classes without the constraints observed in the aforementioned methodologies. To the best of our knowledge, Diffusion models have not yet been explored in the ZSL domain.
@@ 3 Methodology
Problem Setup. We denote the set of seen images, semantics, and corresponding class labels as $\{x^{{seen}},s^{{seen}},y^{{seen}}\}\in\mathcal{D}^{{seen}}$, where $x^{{seen}}$ represents the images, $s^{{seen}}$ the semantics, and $y^{{seen}}$ the class labels for the seen classes. The set of unseen semantics and class labels, denoted as $\{s^{{unseen}},y^{{unseen}}\}\in\mathcal{D}^{{unseen}}$, represents the unseen dataset. During training, the model is trained exclusively on data from the seen set $\mathcal{D}^{{seen}}$, while assuming access to the semantic and label information of the unseen classes in $\mathcal{D}^{{unseen}}$. Importantly, the unseen images $x^{{unseen}}$ are not available during training and are only introduced during the inference phase.
It is important to note that during training, the set of class labels for seen and unseen data do not overlap, i.e., $\mathcal{Y}^{seen}\cap\mathcal{Y}^{unseen}=\emptyset$.
During inference, the challenge is to map an unseen sample image, $x^{unseen}$, to its corresponding unseen label $y^{unseen}$, using a learned function $f:x^{unseen}\rightarrow y^{unseen}$.
The training process involves using paired examples $\{x^{seen},s^{{seen}}\}\in\mathcal{D}^{{seen}}$ to learn a model, $p_{\theta}(s|x^{{seen}})$, that generalizes this mapping function from the visual features $x^{{seen}}$ to their corresponding semantic features $s^{{seen}}$ embedded in a shared semantic space. During the test phase, the semantic distribution of the unseen images $x^{{unseen}}$ is approximated and subsequently classified into their corresponding unseen classes $y^{{unseen}}\in\mathcal{Y}^{{unseen}}$.
@@ 3.1 Diffusion Process
The diffusion process [42] models complex data distributions through a specific Markov chain structure. During training, we start with a clean semantic sample and its corresponding visual features $\{s_{0},x\}\in\mathcal{D}^{seen}$. We incrementally infuse Gaussian noise using a fixed linear Gaussian model, which by using the reparameterization trick [25] can be parameterized by mean $\mu_{t}(s_{t})=\sqrt{\alpha_{t}}s_{t-1}$, and variance $\Sigma(s_{t})=(1-\alpha)\mathbf{I}$ for hierarchical time-steps $t\in[0,T]$. Pre-defining a noise schedule $(\beta_{1},...,\beta_{T})$ allows us to sample from the Markov chain through a fixed forward sequence of time steps as:
$q(s_{t}|s_{t-1})=N(s_{t};\sqrt{\alpha_{t}}s_{t-1},(1-\alpha_{t})\mathbf{I})$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. This forward encoding process has the desired properties of being variance-preserving and completely deterministic, and the final distribution $p(s_{T})$ is a standard Gaussian. Our aim is to learn a reveres process $p(s_{t-1}|s_{t},x)$ which removes the noise of the forward process and can estimate a clean sample $\hat{s}_{0}$ from random noise conditioned on a visual space:
$p_{(}s_{0:T}|x)=p(s_{T})\prod_{t=1}^{T}p(s_{t-1}|s_{t},x)$ (2)
Directly expressing $p$ in closed form is intractable. We instead parameterize $p_{\theta}$ with $\theta$ and approximate the distribution by minimizing the evidence lower bound (ELBO) of the conditional log-likelihood
@@
By conditioning the forward process on the clean example at any given $t$, the diffusion loss can be formulated using Bayes rule as the KL divergence between the analytical truth denoising step $q(s_{t-1}|s_{t},s_{0},\mathbf{x})$ and our approximated denoising step $p_{\theta}(s_{t-1}|s_{t},\mathbf{x})$. The prior loss $D_{KL}[q(s_{T}|s_{0},x)||p(s_{T}|x)]$ can be ignored as it does not contain any trainable parameters and is zero under our assumption. The reconstruction loss $log\ p_{\theta}(s_{0}|s_{1},x)$ is typically minimal and can be safely ignored without affecting the outcome. Therefore, our diffusion objective becomes:
@@
$\mathcal{L}_{Diffusion}=\underbrace{\frac{1}{2\sigma_{q}^{2}(t)}\frac{\bar{\alpha}_{t-1}(1-\alpha_{t})^{2}}{(1-\bar{\alpha}_{t})^{2}}}_{w_{t}}\left[||\hat{s}_{\theta}(s_{t},t,\mathbf{x})-s_{0}||_{2}^{2}\right]$ (6)
The first term is a time-dependent variance weight, where $\sigma_{q}^{t}(t)=\beta_{t}$ [19]. However, empirical research [28] has demonstrated that setting $w_{t}=1$ yields optimal performance and our results from a zero-shot paradigm show similar results.
3.1.1 Cross Hadamard-Addition Embeddings
In the traditional diffusion process, we predict the ground truth of a noisy sample at time $t$. In our approach, however, we further condition this process on visual features. Consequently, the neural network $s_{\theta}$ is trained on the triplet $(s_{t},t,x)$, where $\{s,x\}\in\mathcal{D}^{seen}$. To refine the embeddings for both the conditioning variable $x$ and the time-step schedule $t$, we employ a cross Hadamard-Addition method, which enhances the representation and integration of these features within the network. During the representational mapping stage within the network, we use Hadamard integration for the time-step input, acknowledging that the added noise is entirely deterministic, while integrating the visual condition through addition (refer to the encoding step in Fig. 2(b)). In contrast, during the network’s generative stage, we reverse these roles, applying Hadamard integration for a stronger conditional reconstruction and a more relaxed incorporation of the time-step input (see the decoding step in Fig. 2(b)). We observed that this approach resulted in a closer alignment of the joint probability space with the reconstructed features, leading to improved accuracy.
Time-dependent embedding. To increase the dimension of the time step, $t$, we employ a sinusoidal time embedding $\bar{t}\leftarrow TE(t,d)$:
where $d$ is the embedding dimension and $f_{i}$ are frequencies. The temporal encoding dimension is matched layer-wise with the network to better learn a denoising function. This is crucial for accurately reconstructing $s\in S_{\text{seen}}$ and preparing for generalization to $S_{\text{unseen}}$.
Visual-dependent embedding. We implement a Transformer encoder [43] to extract visual features for conditioning. These visual features are integrated into the network at each layer, with the Transformer trained concurrently. To align the denoising feature dimensions, we map the multi-head attention outputs from the visual space to each intermediate feature using a Hadamard product in the decoder and matrix addition in the encoder of our denoising model (see Fig. 2(b)).
3.2 Noise Loss
The diffusion loss in Eq. (4) can also be interpreted as estimating the source noise $\hat{\epsilon}_{t}$, rather than directly predicting the clean sample $\hat{s}_{0}$. By applying the reparameterization trick [34], we can express the relationship between a clean and an arbitrarily noised sample as:
@@ 3.2 Noise Loss
$\mu_{\theta}(s_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}s_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha_{t}}}\sqrt{\alpha_{t}}}\hat{\epsilon}_{t}$ (9)
@@
$\mathcal{L}_{noise}=\underbrace{\frac{1}{2\sigma_{q}^{2}(t)}\frac{(1-\alpha_{t})^{2}}{(1-\bar{\alpha}_{t})^{2}\alpha_{t}}}_{w^{\prime}_{t}}[||\epsilon_{0}-(\hat{s}_{\theta}(s_{t},t,x)-s_{0})||_{2}^{2}]$ (10)
@@
These two complementary formulations of the denoising transition mean correspond to an equivalent optimization problem (Eq. 5). While they introduce additional complexity to the optimization process, necessitating more sophisticated strategies and careful hyperparameter tuning to achieve convergence, we observe significant improvements in feature generation. We believe that this loss function acts as a regularizer during optimization, enhancing the model’s ability to generalize to unseen data. By optimizing from multiple perspectives, the model generates richer and more robust representations.
3.3 Classification loss
Unlike other generative models, like flow-based models and GANs, Diffusion models have no natural property to decrease the intra-class variance from the noise input [20]. Previous work in classifier- and classifier-free guidance in score-based Diffusion models [20] involves modifying the score function with the gradient of the log-likelihood of a separate classifier model $-\eta\nabla_{\vartheta}logp_{\vartheta}(y|s_{t})$. Our classification objective is to steer our optimization problem of the inferred distribution through manifold regularization [47], leveraging a classifier
$\operatorname*{arg\,max}_{\theta}\mathbb{E}_{y\in Y_{seen}}[log\ p_{\vartheta}(\hat{s}_{0:T}|y)]$ (11)
@@ 3.3 Classification loss
$\mathcal{L}_{cls}=-\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{c}y_{ij}\ logp_{\vartheta}(\hat{y}_{ij}|s_{i})$ (12)
@@
where $n$ is the number of samples, $y_{ij}$ and $\hat{y}_{ij}$ is the true and predicted label of for class $j$ of the $i$-th instance.
3.4 Model Architecture

### v2

@@ Abstract
In Generalized Zero-Shot Learning (GZSL), we aim to recognize both seen and unseen categories using a model trained only on seen categories. In computer vision, this translates into a classification problem, where knowledge from seen categories is transferred to unseen ones by exploiting the relationships between visual features and available semantic information. However, learning this joint distribution is costly and requires one-to-one alignment with corresponding semantic information. We present a reversed conditional diffusion-based model (RevCD) that mitigates this issue by estimating the semantic density conditioned on visual inputs. Our RevCD model consists of a cross Hadamard-addition embedding of a sinusoidal time schedule, and a multi-headed visual transformer for attention-guided embeddings. The proposed approach introduces two key innovations. First, we apply diffusion models to zero-shot learning, a novel approach that exploits their strengths in capturing data complexity. Second, we reverse the process by approximating the semantic densities based on visual data, made possible through the classifier-free guidance of diffusion models. Empirical results demonstrate that RevCD achieves competitive performance compared to state-of-the-art generative methods on standard GZSL benchmarks. The complete code will be available on GitHub.
Keywords:
Zero-shot Learning Transfer Learning Diffusion Model
@@ 1 Introduction
Zero-shot learning (ZSL) represents a state-of-the-art advancement in machine learning transferability and computer vision classification. By pushing the boundaries of knowledge extraction, ZSL enable ML models to expand without costly retraining. This learning paradigm is particularly crucial as it addresses the inherent limitation of traditional machine learning models that require prior access to expensive datasets. ZSL leverages auxiliary knowledge, allowing models to explore unobserved events, edge cases, or new compositions without any additional training. Traditional approaches in ZSL focused on aligning attributes directly with object categories [25], while deep learning’s potential to create a joint embedding space of visual and semantic features [8, 34, 7, 12, 40] has rendered this approach obsolete. The shift toward latent-based methods highlights the importance of embedding space techniques because of their ability to decode and infer complex data distributions. This offers a promising resolution to the two main challenges of ZSL: the semantic gap and limited generalization ability [44].
@@
Our contribution introduces a Diffusion-based generative model, a notable innovation in ZSL. Distinct from conventional models that predominantly rely on attribute matching or embedding strategies, our RevCD model utilizes a diffusion process to model the data distributions iteratively, see Fig. 1. This augments the model’s capability to manage class variability and enhances its generalization capacity. Such control over the semantic space is required for overcoming the challenges of bias and hubness commonly encountered in ZSL methodologies [26].
@@ 2 Related works
Advancements in likelihood-based models have been central to the progress of zero-shot learning [36]. By framing the learning process as a maximum likelihood estimation problem, these methods effectively model data distributions, allowing for robust generalization across both seen and unseen classes. This section categorizes ZSL approaches according to the foundational models employed for approximation of the data, including Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Hybrid models. Additionally, it highlights the role of attention mechanisms and embedding strategies in enhancing these models’ performance.
VAE-based. Variational Autoencoders [24] play a crucial role in ZSL due to their probabilistic framework for modeling latent spaces. Their adaptability in synthesizing unseen class prototypes, as demonstrated by [38, 5], and [22], underscores their versatility in ZSL applications. For instance, [18] incorporates a semantic-guided approach within a VAE framework, while [46] employs a decoupling strategy to enhance performance. A significant advantage of VAEs lies in their capacity to explicitly approximate data density. However, a key limitation is their tendency to generate blurry or overly smoothed features due to posterior collapse [32], which can obscure essential class-specific details needed to distinguish between unseen classes.
GAN-based. Generative Adversarial Networks [14] offer a powerful and dynamic framework for feature synthesis. GANs have been successfully adapted to generate features for unseen classes [13, 21, 15, 52]. GANs´ ability to produce sharp and realistic features through implicit density estimation makes them particularly effective for capturing fine-grained details. However, they are also prone to challenges such as training instability and mode collapse [4], which can result in a lack of diversity in the generated features. This limitation may hinder the model’s ability to accurately represent the full spectrum of unseen classes.
Hybrids. Hybrid models in ZSL leverage multiple architectures to enhance performance. The majority of hybrid frameworks incorporate sequential modules, as seen in [16, 10, 30]. Employing a VAE to learn an embedding function that constrains the semantic or visual space allows for greater control over the generation of synthesized features. Nevertheless, the complexity and computational overhead of combining multiple models can pose challenges, especially in terms of model interpretability and scalability.
@@
Our contribution introduces a reversed Diffusion-based model (RevCD) for zero-shot inference. Diffusion models have been applied to improve accuracy as generative classifiers [6, 3, 39], and their capacity to generate synthetic data has enabled the classification of unseen compositions [9, 27]. However, existing applications are limited by pre-training on prompt categories. Their implementation in a pure zero-shot setting is still absent. We address these limitations by leveraging the reversed process for generating conditioned semantic embeddings, aiming for effective generalization to unseen classes without the constraints observed in the aforementioned methodologies. To the best of our knowledge, Diffusion models have not yet been explored within the ZSL domain.
@@ 3 Methodology
3.0.1 Problem Setup.
We denote the set of seen images, semantics, and corresponding class labels as $\{x^{{seen}},s^{{seen}},y^{{seen}}\}\in\mathcal{D}^{{seen}}$, where $x^{{seen}}$ represents the images, $s^{{seen}}$ the semantics, and $y^{{seen}}$ the class labels for the seen classes. The set of unseen semantics and class labels, denoted as $\{s^{{unseen}},y^{{unseen}}\}\in\mathcal{D}^{{unseen}}$, represents the unseen dataset. During training, the model is trained exclusively on data from the seen set $\mathcal{D}^{{seen}}$, while assuming access to the semantic and label information of the unseen classes in $\mathcal{D}^{{unseen}}$. Importantly, the unseen images, $x^{{unseen}}$, are not available during training and are only introduced during the inference phase. It is important to note that during training, the set of class labels for seen and unseen data does not overlap, i.e., $\mathcal{Y}^{seen}\cap\mathcal{Y}^{unseen}=\emptyset$. This ensures there is no direct overlap between seen and unseen classes. During inference, the challenge is to map an unseen sample image, $x^{unseen}$, to its corresponding unseen label $y^{unseen}$, using a learned function $f:x^{unseen}\rightarrow y^{unseen}$. The training process involves using paired examples $\{x^{seen},s^{{seen}}\}\in\mathcal{D}^{{seen}}$ to learn a joint model, $p_{\theta}(s,x)$, that can be used to estimate the conditional probability distribution function for the semantic space $s^{{seen}}$ given their corresponding visual features $x^{{seen}}$ as $p_{\theta}(s|x^{seen})$. During the test phase, the semantic distribution of the unseen images $x^{{unseen}}$ is approximated and subsequently classified into their corresponding unseen classes $y^{{unseen}}\in\mathcal{Y}^{{unseen}}$.
@@ 3.1 Diffusion Process
The diffusion process [41] models complex data distributions through a specific Markov chain structure. During training, we start with time-step 00 represented as an ordinary, clean, sample from the semantic space $s_{0}\sim q(S)$. This is paired with its corresponding visual features $\{s_{0},x\}\in\mathcal{D}^{seen}$. We incrementally infuse Gaussian noise using a fixed linear Gaussian model $s_{1},\dots,s_{T}$, for $T$ steps. By employing the reparameterization trick [24], we can parameterize the mean $\mu_{t}(s_{t})=\sqrt{\alpha_{t}}s_{t-1}$, and variance $\Sigma(s_{t})=(1-\alpha_{t})\mathbf{I}$ for hierarchical time-steps $t\in[0,T]$. Pre-defining the noise schedule $(\beta_{1},\dots,\beta_{T})$ allows us to sample from the Markov chain through a fixed forward sequence of time steps as:
$q(s_{t}|s_{t-1})=N(s_{t};\sqrt{\alpha_{t}}s_{t-1},(1-\alpha_{t})\mathbb{I})$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. This forward encoding process has the desired properties of being variance-preserving and completely deterministic, and the final distribution $p(s_{T})$ is a standard Gaussian. Our aim is to learn a reverse diffusion kernel $p(s_{t-1}|s_{t},x)$ which removes the noise of the forward process and can estimate a clean sample, $\hat{s}_{0}$, from random noise conditioned on a visual space:
$p(s_{0:T}|x)=p(s_{T})\prod_{t=1}^{T}p(s_{t-1}|s_{t},x)$ (2)
Directly expressing $p(\cdot)$ in closed form is intractable. We instead parameterize $p_{\theta}$ with $\theta$ and approximate the distribution by minimizing the evidence lower bound (ELBO) of the conditional log-likelihood
@@
By conditioning the forward process on the clean example at any given $t$, the diffusion loss can be formulated using Bayes´ rule as the KL divergence between the ground-truth analytical denoising step $q(s_{t-1}|s_{t},s_{0},\mathbb{x})$ and our approximated denoising step $p_{\theta}(s_{t-1}|s_{t},\mathbb{x})$. The prior loss $D_{KL}[q(s_{T}|s_{0},x)||p(s_{T}|x)]$ can be ignored as it does not contain any trainable parameters and is zero under our assumption. The reconstruction loss $\log p_{\theta}(s_{0}|s_{1},x)$ is typically minimal and can be safely ignored without affecting the outcome. Therefore, our diffusion objective becomes:
@@
which boils down to learning a neural network, $s_{\theta}$, to predict the semantic space $\hat{s}_{0}$ from noise at time $t$, conditioned on an image $x$. This network can be optimized using stochastic samples of $t$ from a uniform distribution $t\sim\mathcal{U}[1,T]$.
Given our case, where we can set the variances to match exactly, the KL divergence in Eq. (5) can be reduced to a minimization of the difference between the mean of the two distributions [11].
$\operatorname*{arg\,min}_{\theta}D_{KL}(q(s_{t-1}|s_{t},s_{0})||p_{\theta}(s_{t-1}|s_{t},x))$ (6)
$=\operatorname*{arg\,min}_{\theta}\frac{1}{2\sigma_{q}^{2}(t)}[||\mu_{\theta}(\cdot)-\mu_{q}(\cdot)||^{2}_{2}]$ (7)
where $\mu_{q}$ is the ground truth transition mean predetermined in the forward process and $\mu_{\theta}$ is the mean of our learned diffusion kernel.
3.1.1 Reconstruction loss.
Using Bayes’ rule and the Markov assumption, $q(s_{t}|s_{t-1},s_{0})=q(s_{t}|s_{t-1})$ [31], we can show that:
$q(s_{t-1}|s_{t},s_{0})\propto\mathcal{N}(s_{t-1};\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})s_{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})s_{0}}{1-\bar{\alpha}_{t}},\underbrace{\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{I}}_{\sigma_{q}^{2}(t)})$ (8)
@@
$\mu_{q}(\mathbf{s}_{t};s_{0})=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})s_{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})s_{0}}{1-\bar{\alpha}_{t}}$ (9)
$\mathcal{L}_{reconstruction}=\underbrace{\frac{1}{2\sigma_{q}^{2}(t)}\frac{\bar{\alpha}_{t-1}(1-\alpha_{t})^{2}}{(1-\bar{\alpha}_{t})^{2}}}_{w_{t}}\left[||s_{0}-s_{\theta}(s_{t},t,x)||_{2}^{2}\right]$ (10)
Here, we replace $\mu_{q}(\cdot)=s_{0}$ with $\mu_{\theta}(\cdot)=s_{\theta}(s_{t},t,x)$ to estimate the reversed diffusion kernel. The first term, $w_{t}$, is a time-dependent variance weight, where $\sigma_{q}^{t}(t)=\beta_{t}$ [19]. However, empirical research [27] has demonstrated that setting $w_{t}=1$ yields optimal performance. Our experiments in the zero-shot paradigm show consistent outcomes.
3.1.2 Noise loss.
The diffusion loss in Eq. (7) can also be interpreted as estimating the source noise added $\hat{\epsilon}_{t}$, rather than directly predicting the clean sample $\hat{s}_{0}$. By applying the reparameterization trick [33], we can express the relationship between a clean and an arbitrarily noised sample as:
@@
$\mu_{q}(s_{t};s_{0})=\frac{1}{\sqrt{\alpha_{t}}}s_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}\sqrt{\alpha_{t}}}\epsilon_{0}$ (12)
@@
$\mathcal{L}_{noise}=\underbrace{\frac{1}{2\sigma_{q}^{2}(t)}\frac{(1-\alpha_{t})^{2}}{(1-\bar{\alpha}_{t})^{2}\alpha_{t}}}_{w^{\prime}_{t}}[||\epsilon_{0}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}s_{t}}s_{\theta}(s_{t},t,x)||_{2}^{2}]$ (13)
@@
These two complementary formulations of the denoising transition mean correspond to an equivalent optimization problem (Eq. 5). Although these formulations introduce additional complexity to the optimization process, necessitating more sophisticated strategies and careful hyperparameter tuning to achieve convergence, we observe significant improvements in density estimation. We believe that this loss function acts as a regularizer during optimization, since $w^{\prime}_{t}\neq w_{t}$, enhancing the model’s ability to navigate the loss landscape and generalize to unseen data. By optimizing from multiple perspectives, the model generates richer and more robust representations.
3.1.3 Classification loss.
Unlike other generative models such as flow-based models and GANs, Diffusion models have no natural property to decrease the intra-class variance from the noise input [20]. Previous work in classifier- and classifier-free guidance in score-based Diffusion models [20] involves modifying the score function with the gradient of the log-likelihood of a separate classifier model $\vartheta$, $-\nabla_{\vartheta}\log p_{\vartheta}(y|s_{t})$. Our classification objective is to steer our optimization problem of the inferred distribution through manifold regularization [48], leveraging a classifier:
$\operatorname*{arg\,max}_{\vartheta}\mathbb{E}_{y\in Y_{seen}}[\log\ p_{\vartheta}(\hat{s}_{0:T}|y)]$ (14)
@@
$\mathcal{L}_{classification}=-\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{c}y_{ij}\log p_{\vartheta}(\hat{y}_{ij}|s_{i})$ (15)
@@
where $n$ is the number of samples and $y_{ij}$ and $\hat{y}_{ij}$ are the true and predicted label for class $j$ of the $i$-th instance.
3.2 Training Objective
Our main idea focuses on directly modeling the semantic posterior using variational inference rooted in Eq. (16). We achieve this by disentangling the posterior estimation into three key components: noise prediction, data reconstruction, and (auxiliary) classification. This decomposition results in a more complex and nuanced loss landscape [28]. Despite the increased complexity, integrating these distinct loss components enhances the model’s generalization capabilities. This is primarily due to the regularization effects inherent in the multi-faceted loss function and the fine-tuning achieved through careful hyperparameter optimization. To enable classifier-free guidance at inference, we adopt a conditional dropout strategy during training, randomly masking the visual condition when computing the noise prediction loss [20]. This allows the model to learn both conditional and unconditional score estimates within a single unified network. Our overall training objective becomes:
$\begin{split}\mathbf{\mathcal{L}_{total}}=\\ \lambda_{1}\mathcal{L}_{rec}+\lambda_{2}\mathcal{L}_{noise}+\lambda_{3}\mathcal{L}_{classification}\end{split}$ (16)
Here, $\lambda_{1}$ and $\lambda_{3}$ serve as a balancing factors between the objectives of reconstruction and classification, while $\lambda_{2}$ acts as a regularization coefficient. Through this, the probability distribution of the samples aligns with the expectation of the generated conditional samples $p(s)\propto\mathbb{E}_{x\sim p(x)}[p_{\theta}(s|x)]$. The implementation details of this loss function during training are provided in Algorithm (1).
3.3 Architectural considerations
3.3.1 Cross Hadamard-Addition Embeddings.
In the traditional diffusion process, we predict the ground truth of a noisy sample at time $t$. In our approach, however, we further condition this process on visual features. Consequently, the neural network $s_{\theta}$ is trained on the triplet $(s_{t},t,x)$, where $\{s,x\}\in\mathcal{D}^{seen}$. To refine the embeddings for both the conditioning variable $x$ and the time-step schedule $t$, we employ a cross Hadamard-Addition method, which enhances the representation and integration of these features within the network. During the representational mapping stage within the network, we use Hadamard integration for the time-step input, acknowledging that the added noise is entirely deterministic, while integrating the visual condition through addition (refer to the encoding step in Fig. 2(b)). In contrast, during the network’s generative stage, we reverse these roles, applying Hadamard integration for a stronger conditional reconstruction and a more relaxed incorporation of the time-step input (see the decoding step in Fig. 2(b)). We observe that this approach results in a closer alignment of the joint probability space, leading to improved accuracy.
3.3.2 Time-dependent embedding.
To increase the dimension of the time step, $t$, we employ a sinusoidal time embedding $\bar{t}\leftarrow TE(t,d)$:
where $d$ is the embedding dimension and $f_{i}$ are frequencies. The temporal encoding dimension is matched to each layer of the network to learn the denoising function, given any timestep $t$.
3.3.3 Visual-dependent embedding.
We implement a Transformer encoder [43] to extract visual features for conditioning. These visual features are integrated into the network at each layer, with the Transformer trained concurrently. To align the denoising feature dimensions, we map the multi-head attention outputs from the visual space to each intermediate feature using a Hadamard product in the decoder and matrix addition in the encoder of our denoising model (see Fig. 2(b)).
3.3.4 Pre-conditioning.
We also perform an affine transformation of our semantic space $s\in[0,1]$ to resemble a zero-mean Gaussian $s\prime=2\times s-1\in[-1,1]$. This increases the dynamic range leading to better gradient flow and stabilizes convergence as the variance $Var(s_{0})\ll 1$ skews the signal-to-noise ratio when scaled by the noising schedule $\bar{\alpha}_{t}$.
3.4 Model design